# Sparse Ellipsoidal Radial Basis Function Network for Point Cloud Surface Representation

### v2

@@ Abstract
Point cloud surface representation is a fundamental problem in computer graphics and vision. This paper presents a machine learning approach for approximating the signed distance function (SDF) of a point cloud using a sparse ellipsoidal radial basis function network, enabling a compact and accurate surface representation. Given the SDF values defined on the grid points constructed from the point cloud, our method approximates the SDF accurately with as few ellipsoidal radial basis functions (ERBFs) as possible, i.e., represents the SDF of a point cloud by sparse ERBFs. To balance sparsity and approximation precision, a dynamic multi-objective optimization strategy is introduced, which adaptively adds the regularization terms and jointly optimizes the weights, centers, shapes, and orientations of ERBFs. To improve computational efficiency, a nearest-neighbor-based data structure is employed, restricting function calculations to points near each Gaussian kernel center. The computations for each kernel are further parallelized on CUDA, which significantly improves the optimization speed. Additionally, a hierarchical octree-based refinement strategy is designed for training. Specifically, the initialization and optimization of network parameters are conducted using coarse grid points in the octree lattice structure. Subsequently, fine lattice points are progressively incorporated to accelerate model convergence and enhance training efficiency. Extensive experiments on multiple benchmark datasets demonstrate that our method outperforms previous sparse representation approaches in terms of accuracy, robustness, and computational efficiency. The corresponding executable program is publicly available at https://github.com/lianbobo/SE-RBFNet.git.
@@
Surface reconstruction, signed distance function, sparse optimization, radial basis function, multi-objective optimization.
journal: Nuclear Physics B
@@
SE-RBFNet is proposed, where the point cloud surface representation problem is reformulated as a nonlinear sparse optimization problem.
@@
A coarse-to-fine hierarchical optimization strategy based on octree lattice structure is developed, and a nearest-neighbor filtering mechanism is incorporated to improve efficiency and reduce model complexity.
An adaptive basis function addition mechanism is designed, in which new kernel centers are iteratively selected from high-error regions and ellipsoidal parameters are dynamically updated, enabling improved accuracy and faster convergence.
@@ 1 Introduction
Point cloud surface representation is a fundamental technique in computer graphics, computer vision, and 3D modeling [1]. Among the various representation paradigms, signed distance functions (SDFs) have gained widespread adoption due to their ability to implicitly and compactly encode complex surface geometries as the zero-level set of a continuous function [2, 3]. This representation not only facilitates surface reconstruction but also enables efficient storage, editing, and physical simulation. Despite their advantages, SDF-based methods face an inherent trade-off between representation fidelity and computational efficiency. High-fidelity SDF approximations typically require dense sampling or a large number of basis functions, which incurs significant memory and computation costs. On the other hand, sparse approximations often fail to capture fine geometric detail, especially in the presence of anisotropic features.
Classical implicit surface methods, including radial basis function (RBF) interpolation [4, 5, 6, 7, 8] and Poisson surface reconstruction [9, 10, 11], represent surfaces as the zero-level set of a fitted continuous function. These techniques are robust to noise and produce smooth, watertight surfaces, yet they suffer from two major limitations: (1) conventional isotropic RBFs inefficiently model directional or elongated surface structures, often requiring excessive basis functions to fit complex geometry; (2) the dense solvers typically employed in these frameworks scale poorly with increasing data size, limiting their practical usability in large-scale scenarios. More recently, learning-based methods have introduced neural networks for SDF representation [12, 13, 14, 15]. While these approaches improve generalization and surface quality, they rely on computationally expensive training pipelines and lack explicit control over the sparsity and structure of the learned representation.
@@
In this paper, motivated by the success of ellipsoidal radial basis functions (ERBFs) in molecular surface representation [16] and their use in Gaussian splatting [17], we propose the Sparse Ellipsoidal Radial Basis Function Network (SE-RBFNet), a fast and efficient machine learning approach for representing point cloud surfaces by approximating SDFs. Our method shares similarities with SparseRBF [1], which also employs RBFs for the sparse representation of point cloud surfaces. There are two important differences between the two methods. First, SparseRBF typically utilizes spherical basis functions, optimizing only the weight coefficients. This design aims to transform the problem into a linear Least Absolute Shrinkage and Selection Operator (LASSO) [18] problem, thereby accelerating the optimization process but lacking the flexibility to accurately fit complex surfaces. In contrast, SE-RBFNet adopts ERBFs and performs joint optimization of the center positions, the rotation angles along the principal axes, axis lengths, and the weight coefficients. This is a high dimensional nonlinear nonconvex optimization problem. The proposed comprehensive optimization scheme significantly enhances the model’s capacity to represent complex geometric features while requiring a much smaller number of basis functions. The second difference lies in the initialization of the center points. SparseRBF positions the center points along the central axis of the point cloud object, whereas SE-RBFNet uses an inscribed sphere strategy to initialize the center points, improving the adaptability to various surface structures.
@@
In addition, we initially applied radial basis networks for sparse representation in the image domain [19], and later extended this approach to 3D point cloud surfaces [20]. Compared to our earlier work, SE-RBFNet incorporates a multi-objective optimization strategy [21], which adaptively balances L2 error term and L1 regularization term of the basis coefficients to achieve an optimal trade-off between representation accuracy and sparsity. To further improve computational efficiency and robustness, a nearest-neighbor search method is used to identify points near the center of each Gaussian kernel during both the forward and backward steps of optimization, effectively reducing computational complexity. The optimization process is further accelerated via CUDA-based parallel computation. Moreover, we design a hierarchical optimization strategy based on octree structures [22]. Specifically, the initialization and optimization of ERBF parameters are first conducted using coarse grid points in the octree lattice structure. As the optimization progresses, finer lattice points are progressively incorporated into the training set. In addition, extreme error points—identified based on the L2 error at grid points—are introduced as new kernel centers. This hierarchical and iterative process not only improves parameter accuracy but also effectively accelerates the convergence of the training process.
@@
    SE-RBFNet is proposed, where the point cloud surface representation problem is reformulated as a nonlinear sparse optimization problem (see Section 3.2).
@@
    A coarse-to-fine hierarchical optimization strategy based on octree lattice structure is developed, and a nearest-neighbor filtering mechanism is incorporated to improve efficiency and reduce model complexity (see Section 3.4.2).
@@
    An adaptive basis function addition mechanism is designed, in which new kernel centers are iteratively selected from high-error regions and ellipsoidal parameters are dynamically updated, enabling improved accuracy and faster convergence. (see Section 3.4.3).
@@ 2 Related Work
Surface representation from point clouds has been a longstanding topic of research, with implicit representations being particularly effective in capturing both local geometry and global surface topology. In this section, we briefly review classical and learning-based implicit methods.
@@ 2.1 Classical Implicit Representations
Classical implicit methods typically represent the surface as the zero-level set of a continuous function, often a SDF. This strategy enables flexible surface representation and efficient reconstruction via methods like the Marching Cubes algorithm [23].
@@
Early work by Curless and Levoy [24] introduced volumetric SDF fusion, laying the foundation for subsequent approaches. Calakli and Taubin [2] further emphasized that the target implicit function should approximate the true SDF, rather than merely filling volume. Pan and Skala [25] proposed a continuous global optimization framework that minimizes a variational energy combining data fidelity and smoothness, enabling robust surface reconstruction from noisy or incomplete oriented point clouds. In a subsequent work, Pan and Skala [26] improved upon their previous approach by incorporating higher-order derivative regularization into the optimization, further enhancing surface smoothness. Poisson surface reconstruction (PSR) and its variants [9, 10, 27] transform reconstruction into solving a spatial Poisson equation, generating smooth and globally consistent surfaces. Building upon the PSR, Hou et al. [28] proposed an iterative PSR method that estimates normals from reconstructed surfaces in each iteration, progressively enhancing surface quality.
@@
RBFs are another popular choice for implicit surface modeling. These methods place an RBF at each point and optimize the corresponding weights to fit the surface. Carr et al. [4] proposed multi-harmonic RBFs for smooth surface interpolation, and introduced a greedy algorithm, which iteratively added centers associated with large residuals, reducing the total number of basis functions. Ohtake et al. [7] decomposed the global approximation problem into overlapping local subproblems, each solved via least-squares RBF fitting. This significantly reduces the number of required basis functions, enabling sparse yet efficient surface representation. Samozino et al. [8] proposed selecting RBF centers directly from the Voronoi vertices computed from the input point cloud, rather than placing them on the surface or its offset. This strategy yields a more uniform spatial distribution of centers. However, since the fitting process treats both surface and off-surface points equally, approximation errors near the true surface can increase. Further developments on RBF-based surface modeling have explored anisotropic kernels [29], compactly supported RBFs [30, 31], orthogonal least squares (OLS) center selection [32], and Hermite RBFs, which directly incorporate normal information [33, 34, 35]. In large-scale contexts such as geographic surface modeling, space-partitioned and comparative studies [36] have demonstrated both the potential and the limitations of conventional RBFs in handling massive point sets efficiently. More recently, RBF-based surface reconstruction has been further extended through sparse center selection [20], partitioned formulations [37], and novel interpolation schemes [38], broadening its applicability across different surface fitting scenarios.
Beyond RBFs, other implicit representations include Implicit Moving Least Squares (IMLS) [39, 40], Fourier bases [41], and Gaussian formulations [42, 43]. These approachs further highlight the versatility and robustness of implicit methods in surface representation. Another important family is implicit moving least squares (IMLS), first proposed by Shen et al. [39]. IMLS fits a local weighted least squares plane around each query point, constructing an implicit function whose zero-level set defines the surface. To improve scalability and adaptivity for large-scale point clouds, subsequent works have introduced enhancements such as adaptive neighborhood selection [44], provably correct projections [45], floating scale support [40], and deep learning-based acceleration [46].
@@ 2.2 Neural Implicit Representations
Recent methods leverage neural networks to learn implicit shape representations. Mescheder et al. [47] introduced Occupancy Networks, parameterizing implicit functions via neural networks and latent codes. Park et al. [3] proposed DeepSDF, which represents shapes as continuous SDFs modeled by multilayer perceptrons (MLPs). Peng et al. [48] further extended this idea by leveraging 3D convolutional neural networks (CNNs) to capture shape priors.
While these methods achieve high reconstruction quality, they often rely on curated datasets and struggle to generalize to unseen shapes. To address this, Gropp et al. [14] incorporated geometric regularization to enhance generalization. Ben-Shabat et al. [49] introduced soft second-order constraints, and Wang et al. [50] minimized the Hessian determinant near surfaces to suppress artifacts.
In contrast to global shape models, Points2Surf [13] and PPSurf [15] focus on local, patch-based learning with self-supervision, enhancing robustness to noise, sparsity, and varying sampling density. Ma et al. [51] further relaxed supervision requirements by proposing an unsupervised learning framework that estimates surface geometry directly from point clouds without ground-truth SDFs.
These studies highlight the importance of SDFs as a fundamental representation of surfaces, which motivates our adoption of sparse ellipsoidal RBF networks for efficient SDF approximation.
Our work follows the learning-based SDF approximation paradigm by adopting a sparse ellipsoidal RBF network to approximate the SDF, thereby representing the implicit surface as a combination of ellipsoids. Compared to traditional RBF-based approaches, our method formulates a nonlinear optimization problem by optimizing both kernel centers and shapes. It incorporates ellipsoidal anisotropy, sparse center selection, and hierarchical optimization to enhance both efficiency and accuracy. Rather than directly focusing on surface reconstruction, our method aims to construct a compact and sparse representation of the SDF, from which the point cloud surface can be naturally recovered as the level set.
@@ 3 Method
In this section, we will describe in detail the principle of SE-RBFNet for point cloud surface representation. ERBFs are used as nodes in each hidden layer of the neural network. Sparsity is introduced to optimize the representation of the point cloud surface using fewer hidden functions. A detailed process flow of SE-RBFNet is shown in Figure 1.
Figure 1: The workflow and results of the SE-RBFNet. (a) shows the construction of the octree, where the signed distance of each grid point of the octree is calculated and used for model training; (b) shows the initial structure of SE-RBFNet and the structure after sparse optimization. SE-RBFNet takes SDF values on the octree grids and the point cloud as input, and outputs the optimized parameters of ERBFs; (c) shows the reconstructed surfaces obtained from the original SDF values on octree grids (in green) and from the optimized SE-RBFNet approximated SDF values on octree grids (in purple); (d) shows the process of extracting the initial ERBF centers using maximum inscribed spheres. The left part illustrates the interior (in green) and exterior (in tomato red) octree grid points. The right part shows the computed maximum inscribed spheres (in light blue) derived from the interior points; (e) shows that the number of optimized ERBF bases is dramatically reduced while the surface shape is preserved.
@@ 3.1 RBF Networks
Despite their theoretical expressiveness, traditional RBF networks typically employ isotropic Gaussian basis functions with fixed shape, making them less effective in capturing anisotropic structures or highly detailed variations in complex surfaces. In the context of 3D implicit surface representation, this limitation becomes especially pronounced when the input data exhibits uneven curvature, fine-grained features, or elongated geometries that cannot be efficiently modeled by spherical support regions. Moreover, the approximation quality of RBFs is highly sensitive to the choice of kernel shape parameters. Skala et al. [55] showed that assigning individual shape parameters to each basis improves flexibility, but also introduces many local optima, making robust estimation challenging.
To address this limitation, we use an extension of the classical RBF framework: the SE-RBFNet. By replacing spherical kernels with learnable ERBFs [56], our model gains the ability to adapt its receptive field shape, orientation, and scale to better align with the underlying geometry. This anisotropic flexibility not only enhances the fidelity of surface representation but also improves parameter efficiency, as fewer ellipsoidal kernels may be required to represent complex structures compared to their isotropic counterparts.
Furthermore, by integrating this ellipsoidal formulation into a sparse optimization framework, SE-RBFNet retains the universal approximation property of RBF networks while significantly improving adaptability to geometric variations in 3D data.

### v4

# Approximating Signed Distance Fields With Sparse Ellipsoidal Radial Basis Function Networks: A Dynamic Multi-Objective Optimization Strategy

@@ Abstract
Accurate and compact representation of signed distance functions (SDFs) of implicit surfaces is crucial for efficient storage, computation, and downstream processing of 3D geometry. In this work, we propose a general learning method for approximating precomputed SDF fields of implicit surfaces by a relatively small number of ellipsoidal radial basis functions (ERBFs). The SDF values could be computed from various sources, including point clouds, triangle meshes, analytical expressions, pretrained neural networks, etc. Given SDF values on spatial grid points, our method approximates the SDF using as few ERBFs as possible, achieving a compact representation while preserving the geometric shape of the corresponding implicit surface. To balance sparsity and approximation precision, we introduce a dynamic multi-objective optimization strategy, which adaptively incorporates regularization to enforce sparsity and jointly optimizes the weights, centers, shapes, and orientations of the ERBFs. For computational efficiency, a nearest-neighbor-based data structure restricts computations to points near each kernel center, and CUDA-based parallelism further accelerates the optimization. Furthermore, a hierarchical refinement strategy based on SDF spatial grid points progressively incorporates coarse-to-fine samples for parameter initialization and optimization, improving convergence and training efficiency. Extensive experiments on multiple benchmark datasets demonstrate that our method can represent SDF fields with significantly fewer parameters than existing sparse implicit representation approaches, achieving better accuracy, robustness, and computational efficiency. The corresponding executable program is publicly available at https://github.com/lianbobo/SE-RBFNet.git.
@@
Signed distance function, implicit surface representation, sparse optimization, radial basis function, multi-objective optimization.
@@
SE-RBFNet is proposed, where the sparse representation of signed distance functions is reformulated as a nonlinear sparse optimization task over the precomputed SDF samples.
@@
A coarse-to-fine hierarchical optimization strategy based on SDF grid points is developed, and a nearest-neighbor filtering mechanism is incorporated to improve efficiency and reduce model complexity.
An adaptive basis function addition mechanism iteratively selects new kernel centers from high-error regions and dynamically updates ellipsoidal parameters, enabling improved accuracy and faster convergence.
@@ 1 Introduction
Accurate and compact representation of signed distance functions (SDFs) is fundamental for efficient storage, computation, and downstream processing of 3D geometry [30]. SDFs have gained widespread adoption due to their ability to implicitly encode complex surface geometries as the zero-level set of a continuous function [4, 45]. This implicit representation facilitates tasks such as surface reconstruction, storage, editing, and physical simulation.
Most implicit surface representation methods, including radial basis function (RBF) interpolation [5, 12, 55, 41] and neural implicit methods [18, 15, 17, 58] operate directly on point clouds or meshes, learning SDFs from raw geometric data. While these methods can achieve high-fidelity implicit surfaces, they often require a large number of parameters to capture fine-scale or anisotropic features. Motivated by the success of ellipsoidal radial basis functions (ERBFs) in molecular surface representation [19] and their use in Gaussian splatting [27], we propose the Sparse Ellipsoidal Radial Basis Function Network (SE-RBFNet), a fast and efficient machine learning approach for sparse representation of precomputed SDFs. Unlike previous methods, SE-RBFNet does not directly rely on raw point clouds. Instead, it takes precomputed SDF samples—which can be generated by existing point-cloud-based or neural implicit methods, analytical SDFs, or triangle mesh evaluation—as input. By approximating these SDF values using a small set of ERBFs, our method produces a sparser representation of the same implicit surface, achieving further compression in parameter count while preserving geometric accuracy near the zero-level set. Our formulation is thus complementary to existing SDF generation techniques, as it allows their outputs to be efficiently sparsified through subsequent approximation.
@@
Our method shares similarities with SparseRBF [30] in its use of RBFs for sparse representation but differs in two key aspects. First, SparseRBF typically utilizes spherical basis functions, optimizing only the weight coefficients. This design aims to transform the problem into a linear Least Absolute Shrinkage and Selection Operator (LASSO) [54] problem, thereby lacking the flexibility to accurately fit complex surfaces. In contrast, SE-RBFNet adopts ERBFs and jointly optimizes basis centers, the rotation angles along the principal axes, axis lengths, and the weight coefficients. This is a high-dimensional nonlinear nonconvex optimization problem. This richer parameterization enables high-accuracy approximation of SDF values, allowing fine-scale and elongated surface regions to be represented using far fewer basis functions. Furthermore, instead of initializing centers along a predefined axis, we employ an inscribed-sphere initialization strategy, improving adaptability to diverse surface shapes.
@@
In addition, we initially applied radial basis networks for sparse representation in the image domain [8], and later extended this approach to 3D surface domains [57]. Compared with our earlier work in [57], SE-RBFNet incorporates a multi-objective optimization strategy [50], which adaptively balances L2 error term and L1 regularization term of the basis coefficients to achieve an optimal trade-off between approximation accuracy and sparsity. To further improve computational efficiency and robustness, a nearest-neighbor search method is used to identify points near the center of each Gaussian kernel during both the forward and backward steps of optimization, effectively reducing computational complexity. The optimization process is further accelerated via CUDA-based parallel computation. Moreover, we design a hierarchical optimization strategy based on SDF grid points. Specifically, ERBF parameters are first initialized and optimized using coarse grid points. As the optimization progresses, finer grid points are progressively incorporated into the training set. This strategy applies to both uniform grids and adaptive grids such as octrees [59]. In addition, extreme error points—identified based on the L2 error at grid points—are introduced as new kernel centers. This coarse-to-fine, iterative process enhances SDF approximation accuracy while accelerating convergence.
@@
    SE-RBFNet is proposed, where the sparse representation of SDFs is reformulated as a nonlinear sparse optimization task over the precomputed SDF samples (see Section 3.2).
@@
    A coarse-to-fine hierarchical optimization strategy based on SDF grid points is developed, and nearest-neighbor filtering is incorporated to improve efficiency and reduce model complexity (see Section 3.4.2).
@@
    An adaptive basis function addition mechanism is designed, in which new kernel centers are iteratively selected from high-error regions and ellipsoidal parameters are dynamically updated, enabling improved accuracy and faster convergence (see Section 3.4.3).
@@ 2 Related Work
Implicit surface representation has been a longstanding topic of research. Many methods aim to directly infer implicit functions—such as SDFs or occupancy fields—from raw geometric data. In contrast, our work focuses on a different problem: efficiently representing a precomputed set of SDF samples in a compact and memory-efficient manner, independent of how these SDF values were obtained. This separation between SDF estimation and representation enables us to focus on efficient encoding while fully leveraging high-quality SDF inputs. In the following, we briefly review classical and learning-based implicit surface representation methods.
@@ 2.1 Classical Implicit Representations
Classical methods typically represent a surface as the zero-level set of a continuous scalar field, often an SDF. This strategy enables flexible shape representation and efficient surface extraction using algorithms such as Marching Cubes [34]. Early work by Curless and Levoy [10] introduced volumetric SDF fusion, laying the foundation for subsequent approaches. Calakli and Taubin [4] further emphasized that the target implicit function should approximate the true SDF, rather than merely filling volume. Pan and Skala [43] proposed a continuous global optimization framework that minimizes a variational energy combining data fidelity and smoothness, enabling robust surface representation from noisy or incomplete oriented point clouds. In a subsequent work, Pan and Skala [44] improved upon their previous approach by incorporating higher-order derivative regularization into the optimization, further enhancing surface smoothness. Poisson surface reconstruction (PSR) and its variants [23, 24, 26, 49] transform reconstruction into solving a spatial Poisson equation, generating smooth and globally consistent surfaces. Building upon the PSR, Hou et al. [20] proposed an iterative PSR method that estimates normals from reconstructed surfaces in each iteration, progressively enhancing surface quality.
@@
RBFs are another popular choice for implicit surface modeling. These methods place an RBF at each point and optimize the corresponding weights to fit the surface. Carr et al. [5] proposed multi-harmonic RBFs for smooth surface interpolation and introduced a greedy algorithm, which iteratively added centers associated with large residuals, reducing the total number of basis functions. Ohtake et al. [41] decomposed the global approximation problem into overlapping local subproblems, each solved via least-squares RBF fitting. This significantly reduces the number of required basis functions, enabling sparse yet efficient surface representation. Samozino et al. [48] proposed selecting RBF centers directly from the Voronoi vertices computed from the input point cloud, rather than placing them on the surface or its offset. This strategy yields a more uniform spatial distribution of centers. However, since the fitting process treats both surface and off-surface points equally, approximation errors near the true surface can increase. Further developments on RBF-based surface modeling have explored anisotropic kernels [6], compactly supported RBFs [56, 42], orthogonal least squares (OLS) center selection [60], and Hermite RBFs [3, 37, 32], which directly incorporate normal information. In large-scale contexts such as geographic surface modeling, space-partitioned and comparative studies [38] have demonstrated both the potential and the limitations of conventional RBFs in handling massive point sets efficiently. More recently, RBF-based surface representation has been further extended through sparse center selection [57], partitioned formulations [13], and novel interpolation schemes [61], broadening its applicability across different surface fitting scenarios.
Beyond RBFs, alternative implicit formulations include Implicit Moving Least Squares [52, 16, 33], Fourier bases [25], and Gaussian-based formulations [35, 31], further illustrating the versatility of implicit representations.
@@ 2.2 Neural Implicit Representations
Recent learning-based approaches parameterize implicit functions using neural networks. Mescheder et al. [40] introduced Occupancy Networks, parameterizing implicit functions via neural networks and latent codes. Park et al. [45] proposed DeepSDF, which represents shapes as continuous SDFs modeled by multilayer perceptrons (MLPs). Extensions include convolutional neural networks (CNNs) for capturing shape priors [47], geometric regularization [17], second-order constraints [2], and Hessian-based smoothness [58]. In contrast to global shape models, Points2Surf [15] and PPSurf [14] focus on local, patch-based learning with self-supervision, enhancing robustness to noise, sparsity, and varying sampling density. Ma et al. [36] further relaxed supervision requirements by proposing an unsupervised learning framework that estimates surface geometry directly from point clouds without relying on ground-truth SDFs.
These studies highlight the importance of SDFs as a fundamental surface representation. Most existing methods—whether classical implicit models or neural implicit models—focus on estimating or learning SDFs directly from raw data, typically requiring a large number of function or network parameters to represent the SDF. Complementary to these approaches, our work targets the sparse representation of precomputed SDF samples, approximating the implicit surface as a combination of ellipsoids. This design enables our method to operate in a data-agnostic manner, leveraging SDFs from any source, including oriented point clouds, triangle meshes, analytically generated SDFs, or SDFs produced by the previously discussed methods. As a general SDF representation framework, SE-RBFNet can efficiently encode a given set of SDF samples with high approximation accuracy and sparsity.
@@ 3 Method
In this section, we describe the SE-RBFNet framework for approximating precomputed SDFs to achieve a compact and sparse representation of implicit surfaces. ERBFs are used as nodes in each hidden layer of the neural network. Sparsity is introduced to represent the surface using fewer RBFs. The overall workflow of SE-RBFNet is illustrated in Figure 1. SDFs derived from an implicit surface are used here as an example. For clarity, we describe the process using an octree grid to provide spatial sampling points for training and subsequent surface extraction, though other sampling strategies, such as uniform grids, could be used as well.
Figure 1: The workflow and results of the SE-RBFNet on SDFs of implicit surfaces. (a) takes the octree structure as an example, where the signed distances of all grid points can be obtained using arbitrary SDF-generation methods and used as training data; (b) shows the initial structure of SE-RBFNet and the structure after sparse optimization. SE-RBFNet takes the SDF values on the octree grids and the sampling points on the implicit surface as input and outputs the optimized parameters of ERBFs; (c) shows the explicit surfaces extracted from the original SDF values on octree grids (in green) and from the SE-RBFNet approximated SDF values on octree grids (in purple); (d) shows the process of extracting the initial ERBF centers using maximum inscribed spheres. The left part illustrates the interior (in green) and exterior (in black) octree grid points. The right part shows the computed maximum inscribed spheres (in light blue); (e) shows that the number of optimized ERBF bases is dramatically reduced while the surface shape is preserved.
@@ 3.1 RBF Networks
Despite their theoretical expressiveness, traditional RBF networks typically employ isotropic Gaussian basis functions with fixed shape, making them less effective in capturing anisotropic structures or highly detailed variations in complex surfaces. In the context of 3D implicit surface representation, this limitation becomes especially pronounced when the input data exhibits uneven curvature, fine-grained features, or elongated geometries that cannot be efficiently modeled by spherical support regions. Moreover, the approximation quality of RBFs is highly sensitive to the choice of kernel shape parameters. Skala et al. [53] showed that assigning individual shape parameters to each basis improves flexibility but also introduces many local optima, making robust estimation challenging.
To overcome these limitations, SE-RBFNet extends the classical RBF framework by replacing spherical kernels with learnable ERBFs [63]. This allows each kernel to adapt its shape, orientation, and scale to better align with the underlying geometry. The resulting anisotropic flexibility enhances approximation fidelity and parameter efficiency, as fewer ellipsoidal kernels are needed to represent complex structures compared with isotropic ones. Furthermore, by integrating ERBFs into a sparse optimization framework, SE-RBFNet retains the universal approximation property of RBF networks while enabling efficient and compact representation of precomputed SDF samples, improving adaptability to geometric variations in 3D surfaces.